OpenStudy (anonymous):

A force of 4 N acts on a body of mass 2 kg for 4 s. Assuming that the body to be initially at rest find : A) its velocity when force stops acting on it. B) The distance covered in 10 s after force starts acting on it @TuringTest @lgbasallote @Lachlan1996

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

@ParthKohli help

OpenStudy (anonymous):

which one are u solving a or b

OpenStudy (anonymous):

actually sorry no

OpenStudy (anonymous):

Come on Lachan come on!

OpenStudy (anonymous):

hahaha sorry, one sec

OpenStudy (anonymous):

a = v= 8m/s

OpenStudy (anonymous):

with b is this assuming that the force stops acting after 4 seconds? and that there is no friction stopping it?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

then

OpenStudy (anonymous):

well then, assuming no friction and that acceleration stops after 4 seconds. you can just use the basic s=vt. =8*10 = 80m

OpenStudy (anonymous):

agree? disagree?

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

BUDDY are u here

OpenStudy (anonymous):

onesec getting help with maths, pm me

OpenStudy (anonymous):

F/m = a = 4/2 = 2 m/s^2 v = at = 2 x 4 = 8 m/s The distance is harder. We have 6 seconds after the force stops acting (at 8 m/s), and the first 4 seconds with an average velocity of 4, so that's 6 x 8 + 4 x 4. Sound right?

OpenStudy (anonymous):

\[F=DeltaP/DeltaT\] P is the linear momentun T is the time \[DeltaP=m(V-V0)\]

OpenStudy (anonymous):

V is the final velocity after 4s V0=0 because of the particle is initially at rest

OpenStudy (anonymous):

\[V=DeltaT F/m\]

OpenStudy (shane_b):

This should be a clearer explanation: \[F=ma==>a=\frac{F}{m}=\frac{4N}{2kg}=2m/s^2\]For answer a: \[V_{f}=V_{i}+at=0m/s + (2m/s^2)(4s) = 8m/s\] For answer b you need to break it into two distances. First, you need the distance traveled while the force acting on (while it was accelerating): \[d=V_it+\frac{1}{2}at^2=(0m/s)(4s)+\frac{1}{2}(2m/s^2)(4s)^2=16m\] Next you need to calculate the distance while it is moving at a constant velocity of 8m/s (which it does for 6 seconds): \[d=vt=(8m/s)(6s)=48m\]Therefore, the total distance will be \[16m+48m=64m\]